two consecutive negative integers have a product of 240.
what are the integers?
n=
n+1=

Answer by checkley77(12844)

N(N+1)=240
N^2+N-240=0
(N+16)(N-15)=0
N+16=0
N=-16 ANS.
-16+1=-15 ANS.
PROOF:
-16*-15=240
240=240